### v1

# MoNE: Replacing Redundant Experts with Lightweight Novices for Structured Pruning of MoE

Abstract
Mixture-of-Experts (MoE) enables efficient scaling of large language models by activating only a subset of experts per input token. However, deploying MoE-based models incurs significant memory overhead due to the need to retain all experts in memory. While structured pruning is promising to reduce memory costs, existing methods often show suboptimal performance and unstable degradation in three dimensions: model architectures, calibration data sources, and calibration sample sizes. This paper proposes Mixture-of-Novices-and-Experts (MoNE), a novel expert pruning method that replaces redundant experts with lightweight novices to achieve effective and robust model compression. MoNE evaluates expert redundancy based on two metrics: access frequency and output variance. Experts exhibiting low usage and stable outputs are pruned and replaced with lightweight novices—unbiased estimations of their original outputs—minimizing performance degradation. Extensive experiments demonstrate that MoNE consistently outperforms baseline methods with minimal accuracy degradation across the three dimensions, confirming its effectiveness and robustness. Notably, it improves the average zero shot accuracy across nine downstream tasks by up to 2.71 under 25% pruning ratio and 3.61 under 50% pruning ratio [^1] The code is available at https://github.com/zxgx/mode-pd..
1 Introduction
Mixture-of-Experts (MoE) has emerged as a powerful architecture for advancing the capabilities of large language models (LLMs) [20, 21, 25]. MoE-based LLMs achieve higher parameter efficiency than vanilla transformer-based LLMs by replacing the MLP module with a set of smaller MLP modules (experts) and sparsely activating partial amount of experts for each input token [17]. Despite its performance benefits, the deployment of MoE-based models often incur additional memory overhead to maintain the non-activated experts in memory, which is valuable but limited for existing accelerators such as GPU and TPU [16].
While diverse structured pruning methods have been proposed to reduce deployment memory costs by removing different model components while minimizing the performance degradation [32, 12, 34], we observe that these approaches often exhibit suboptimal performance and unstable degradation when applied to MoE models. Specifically, we identify three critical dimensions where existing methods fall short: model architectures, calibration data sources and calibration sample sizes, as shown by experiments in Section 5.3. These limitations are evident across two main categories of structured pruning approaches for MoE models: general structured pruning and expert pruning as shown in Figure 1 (a). First, general structured pruning methods that remove model layers (Angular [11]) or weight matrix channels (FLAP [2]) fail to account for the sparse computation scheme of MoE models when evaluating component importance, resulting in inconsistent performance across the aforementioned three dimensions. Second, existing expert pruning methods such as MC-SMoE [18] and RS [12] remove experts from MoE models primarily based on the expert access frequency. However, as shown in Figure 1 (b), this feature alone fails to fully capture the expert redundancy. Besides, these methods lack mechanisms to recover the performance loss caused by pruning.
Figure 1: (a) Different structured pruning methods. (b) Layer-wise normalized expert access frequency and output variance of Deepseek-V2-Lite for three downstream tasks. Experts with high access frequency or output variances are the same across downstream tasks. Expert in blue circles has both high frequency and variance. Expert in red circles only has high variance. Expert in green circles only has high frequency. Similar observations on other models and tasks are in Appendix B.
To improve the effectiveness and robustness of structured pruning for MoE models, this paper proposes a novel expert pruning method, Mixture-of-Novices-and-Experts (MoNE) which replaces redundant experts with a lightweight structure, novice. Specifically, to prune an MoE model with MoNE, it first evaluates the expert redundancy by the access frequency and the output variance for each expert on a calibration dataset. Then, it identifies and removes redundant experts that show low access frequency and stable output activations. Finally, the unbiased estimation of the expert output is employed as the lightweight novice to reduce the computation and memory overhead for the pruned expert. The intuition behind MoNE is that experts with low access frequency contribute less to the final outputs and experts whose outputs have low variance can be replaced with a constant but introduce less discrepancy. Moreover, Figure 1 (b) reveals that experts with less redundancy identified by MoNE exhibit strong consistency across various downstream tasks.
The contribution of this paper is summarized as follows:
  • We propose a novel expert pruning method named MoNE which replaces redundant experts with lightweight novices to compress MoE models.
  • We exploit the expert access frequency and output variance to measure the expert redundancy and employ the unbiased estimation of the expert output to minimize the output discrepancy after pruning, thus achieving effective and robust pruning results.
  • Extensive experiment results demonstrate that MoNE consistently outperforms baseline methods under varying MoE architectures, calibration data sources and calibration sample sizes. Notably, it improves the average zero shot accuracy across nine downstream tasks by up to 2.71 with 25% pruning ratio and 3.61 with 50% pruning ratio.
2 Related Work
Model pruning compresses a model by removing certain redundant model parameters while preserving accuracy. Existing pruning methods generally fall into two categories: unstructured pruning and structured pruning. Unstructured pruning eliminates any model parameter that has minimal impact on model performance. Methods such as SparseGPT [9], Wanda [30], SparseLLM [3] excel in maintaining accuracy while achieving high compression ratios. However, the resulting irregular sparsity patterns hinder efficient representation and execution on hardware accelerators.
In contrast, structured pruning removes certain modules of a model, preserving hardware-friendly structures. Early researches prune redundant transformer layers of a LLM [8, 19, 11]. LLM-Pruner [23] and FLAP [2] remove rows or columns of individual weight matrices. Recent work also proposes to delete components such as attention, MLP or MoE modules within each transformer layer [32, 12]. Minitron [26] and Sheared LLaMA [34] combine different granularity and automatically search for the optimal structures to prune. Despite their versatility, existing structured pruning methods often exhibit inconsistent performance across MoE architectures.
This work focuses on expert pruning, a unique direction of structured pruning for MoE models [17, 20]. Expert pruning targets on deleting individual experts for each layer to compress an MoE model. Previous expert pruning methods either require exhaustive search to identify redundant experts [22], or heavily rely on retraining to recover accuracy due to the suboptimal pruning performance [18, 12]. However, the exhaustive search is not applicable to modern MoE model architectures such as Deepseek [7, 21] or OLMoE [25], as their MoE layer contains 64 experts or even more, yielding a tremendous search space that is intractable. Retraining obscures the advantages of expert pruning over other structured pruning methods.
3 Preliminaries
3.1 Mixture-of-Experts (MoE)
MoE-based LLMs replace the traditional MLP module in the transformer layer with MoE module. Each MoE module consists of a router network $G$ and a set of experts =$\mathcal{E}=\{E_{1},E_{2},\dots,E_{M}\}$, where $M$ is the number of experts and each expert is a smaller MLP. Let $\textbf{x}\in\mathbb{R}^{d}$ be the hidden state of an input token, where $d$ is the hidden size of the model, the output of an MoE module is computed as:
$\mathrm{MoE}(\textbf{x},G,\mathcal{E})=\sum_{E_{i}\in\mathcal{S}_{k,\textbf{x}}}G_{i}(\textbf{x})\cdot E_{i}(\textbf{x})$ (1)
The output of the router network $G(\textbf{x})\in\mathbb{R}^{M}$ represents the routing scores for all experts, and $\mathcal{S}_{k,\textbf{x}}\subseteq\mathcal{E}$ denotes the top $k$ experts with the highest routing scores for input x. The final output of the MoE module is the weighted sum of outputs from the top $k$ experts. While Equation 1 captures the general MoE computation, implementations for $G$ and $E_{i}$ may vary across model architectures [7, 20, 25, 21].
3.2 Expert pruning formulation
Previous studies have revealed that not all experts contribute equally, and pruning less important ones can reduce memory overhead with marginal performance degradation [22, 18, 14]. However, searching for the target experts to prune at the global model perspective falls into a tremendous search space, as the number of experts per transformer layer increases with the evolving of the MoE model architectures [17, 15, 7, 20]. Following the layer-wise pruning scheme [9, 2, 22, 19], our goal of expert pruning is to identify a subset of redundant experts $\mathcal{P}\subseteq\mathcal{E}$ such that we can minimize the output difference after compressing their parameters:
$\underset{\mathcal{P}\subseteq\mathcal{E}}{\min}\|\mathrm{MoE}(\textbf{x},G,\mathcal{E}\setminus\mathcal{P})-\mathrm{MoE}(\textbf{x},G,\mathcal{E})\|_{2}$ (2)
To achieve this goal, the core problem is twofold: (1) find a metric to evaluate the importance of the experts in each layer, so that we can identify the expert subset $\mathcal{P}$, and (2) find an pruning method to compress the parameters of $\mathcal{P}$, so that we can reduce the model size. While $\mathcal{E}\setminus\mathcal{P}$ implies directly removing redundant experts [12, 22], existing methods have also explored expert merging to mitigate the expert redundancy [18].
Figure 2: The overview of MoNE. Given an MoE model, it first exploits a calibration dataset to evaluate the expert access frequency and output variance. Then, the two metrics are fused to get the expert redundancy. Finally, the novices are derived from the averaged outputs for redundant experts.
4 Mixture of Novices and Experts
This section introduces Mixture-of-Novices-and-Experts (MoNE), a novel expert pruning method designed to achieve effective and robust compression for MoE models while minimizing performance degradation. Section 4.1 presents the computational framework of MoNE. Section 4.2 defines the metric to evaluate the redundancy of experts. Section 4.3 explains the pruning process that compresses the redundant experts to lightweight novices. The overview of MoNE is depicted in Figure 2.
4.1 MoNE framework
MoE models are often trained with auxiliary losses to ensure load balance among experts in each layer, enabling each expert to learn certain aspect of knowledge [17, 7, 25]. However, most existing expert pruning methods directly remove experts [12, 18], often leading to inconsistent performance drops across different model architectures or calibration data. MoNE addresses this issue by introducing lightweight structures called novices to replace the pruned experts. A novice is designed to capture the essential knowledge previously held by the removed experts. In contrast to simply removing redundant experts, MoNE compensates for knowledge loss by leveraging novices, thereby preserving the overall performance of the model while maintaining compression efficiency. Specifically, the output of the MoNE is computed as:
$\mathrm{MoNE}(\textbf{x})=\left(\sum_{E_{i}\in\mathcal{S}_{k,\textbf{x}}\setminus\mathcal{P}}G_{i}(\textbf{x})\cdot E_{i}(\textbf{x})\right)+\left(\sum_{E_{i}\in\mathcal{S}_{k,\textbf{x}}\cap\mathcal{P}}G_{i}(\textbf{x})\cdot N_{i}\right)$ (3)
where $\mathcal{S}_{k,\textbf{x}}\setminus\mathcal{P}$ and $\mathcal{S}_{k,\textbf{x}}\cap\mathcal{P}$ denote the preserved and pruned experts among the top $k$ activated experts respectively. $N_{i}\in\mathbb{R}^{d}$ is the novice $i$ that retains the essential knowledge of the pruned expert. Notably, $N_{i}$ is a compressed vector that does not involve any computation with the input token x. As a result, the computation and memory overhead is nearly identical to directly removing experts. Furthermore, replacing experts with novices introduces adaptive computation overhead for different tokens, leading to fewer activated parameters for tokens routed to novices. Nevertheless, empirical results in Section 5.2 demonstrate that MoE models pruned by MoNE maintain more zero shot performance on downstream tasks compared to existing expert pruning methods that only remove experts but keep the same activated parameters.
4.2 Expert redundancy evaluation
To identify the expert subset $\mathcal{P}$, we introduce an expert redundancy score $\phi$ to assess the redundancy of experts. To ensure the pruned experts contribute minimally to the model’s overall performance, the expert redundancy score $\phi$ takes two aspects into consideration: the variance in an expert’s output across a calibration dataset $\mathcal{C}$, and the frequency of an expert selected by the router network $G$.
Variance-based redundancy
As the novices are constant vectors to ensure reduced computation and memory overhead, the outputs of the pruned experts are expected to have low variance across a calibration dataset $\mathcal{C}$. In other words, experts with high output variance should be retained to contribute more discriminative information during inference, whereas experts with low output variance could be compressed into a more efficient representation, i.e., a novice. The second row of Figure 1 (b) visualizes this motivation. Expert outputs exhibit diverse variances, but we can find experts in blue and red circles that maintain high variances across different downstream tasks. Therefore, we introduce a variance-based redundancy $\phi^{var}_{i}$ to measure the output variance for expert $E_{i}$. Concretely, $\phi^{var}_{i}$ is the L2 norm of the unbiased estimation for the output variance:
$\phi_{i}^{var}$ $=\left\|\sqrt{\frac{\sum_{\textbf{x}\in\mathcal{C}}(E_{i}(\textbf{x})-\overline{E_{i}})^{2}\cdot\mathbb{I}(E_{i}\in\mathcal{S}_{k,\textbf{x}})}{\sum_{\textbf{x}\in\mathcal{C}}\mathbb{I}(E_{i}\in\mathcal{S}_{k,\textbf{x}})-1}}\right\|_{2}$ (4)
$\overline{E_{i}}$ $=\frac{\sum_{\textbf{x}\in\mathcal{C}}E_{i}(\textbf{x})\cdot\mathbb{I}(E_{i}\in\mathcal{S}_{k,\textbf{x}})}{\sum_{\textbf{x}\in\mathcal{C}}\mathbb{I}(E_{i}\in\mathcal{S}_{k,\textbf{x}})}$ (5)
where $\mathbb{I}(E_{i}\in\mathcal{S}_{k,\textbf{x}})$ is the indicator function to show whether $E_{i}$ is among top $k$ experts for the input token x of the calibration dataset $\mathcal{C}$.
Frequency-based redundancy
The routing scores and access frequencies of the router network $G$ serve as strong indicators of the overall redundancy of an expert [12, 18]. Intuitively, experts which are rarely selected or consistently assigned lower routing scores are likely to have a minimal impact on the model’s output. As shown in Figure 1 (b), we can identify typical experts in blue and green circles that show consistent high frequency over the three downstream tasks. Notably, the expert in green circles only has high frequency. Therefore, the frequency and variance information can complement the discrepancy ignored by each other. Based on this observation, we define the frequency-based redundancy $\phi_{i}^{freq}$ of the expert $E_{i}$ as the average routing score across a calibration dataset $\mathcal{C}$ of which $E_{i}$ is among the top $k$ selected experts. Formally, the frequency-based redundancy $\phi_{i}^{freq}$ is defined as:
$\phi_{i}^{freq}=\frac{\sum_{\textbf{x}\in\mathcal{C}}G_{i}(\textbf{x})\cdot\mathbb{I}(E_{i}\in\mathcal{S}_{k,\textbf{x}})}{\sum_{\textbf{x}\in\mathcal{C}}\mathbb{I}(E_{i}\in\mathcal{S}_{k,\textbf{x}})}$ (6)
Finally, the two redundancy metrics are fused to obtain the expert redundancy score $\phi$:
$\phi=\phi^{var}\cdot\phi^{freq}$ (7)
A lower expert redundancy score $\phi_{i}$ indicates higher redundancy for expert $E_{i}$, making it a suitable candidate for pruning and replacement with a novice $N_{i}$.
4.3 Expert replacement with novice
After identifying the pruned expert subset $\mathcal{P}$, we need to construct lightweight novices to replace them. According to Equation 2, the general objective for expert pruning is to minimize the discrepancy introduced by the removed expert outputs. Since the output after applying MoNE is formulated as Equation 3, the concrete objective for MoNE can be translated to:
$\underset{E_{i}\in\mathcal{P}}{\min}\|\sum_{\textbf{x}\in\mathcal{C}}(E_{i}(\textbf{x})-N_{i})\|_{2}$ (8)
Because $N_{i}$ is a constant vector, the optimal novice vector $N_{i}$ that best approximates the output of a pruned expert $E_{i}$ can be obtained in a closed form, i.e., $\overline{E_{i}}$ in Equation 5.
To sum up, MoNE uses the unbiased estimations of mean expert outputs to replace experts that have the minimum output variance. As a result, MoNE achieve the goal that effectively and robustly compresses the MoE experts while minimizing performance degradation.
5 Evaluation
5.1 Experiment setup
Base MoE models
To validate the effectiveness and robustness of MoNE, we conducted structured pruning on three open source MoE models with diverse architectures and model scales: OLMoE [25], Moonlight [21] and DeepSeek-V2-Lite [7]. OLMoE has 7B parameters with 1B activated parameters per token. Both Moonlight and Deepseek-V2-Lite have 16B parameters with 3B activated pamaters per token. OLMoE and Moonlight represent SOTA MoE models at their respective scales. We chose the base version of the three models for experiments [^2] OLMoE: https://huggingface.co/allenai/OLMoE-1B-7B-0125, Moonlight: https://huggingface.co/moonshotai/Moonlight-16B-A3B, Deepseek V2 Lite: https://huggingface.co/deepseek-ai/DeepSeek-V2-Lite.
Table 1: Zero shot performance with 100 calibration samples from Zyda2 dataset. Best results are in bold, and the second best are underlined. Green cells indicate results no less than original models.
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Avg.
0% OLMoE 49.23 76.89 70.09 85.00 53.54 44.40 79.76 71.84 68.90 66.63
25% Angular 32.76 61.91 61.71 74.00 23.13 37.60 71.65 53.07 55.09 52.33
FLAP 40.53 67.55 62.69 78.00 41.16 37.80 74.81 61.37 60.93 58.32
MC-SMoE 35.67 54.92 63.49 73.00 29.04 30.60 67.19 55.23 65.75 52.77
RS 25.85 43.01 59.08 74.00 29.63 36.20 66.16 56.68 59.98 50.07
MoNE 42.32 64.81 67.19 85.00 40.13 40.80 78.07 64.62 66.46 61.04
50% Angular 27.22 37.50 53.91 62.00 23.96 26.60 58.27 52.35 51.85 43.74
FLAP 29.18 54.92 62.17 68.00 30.51 29.60 67.57 55.23 56.27 50.39
MC-SMoE 24.49 31.44 59.33 67.00 23.01 26.00 53.92 53.43 53.12 43.53
RS 21.50 28.62 39.45 61.00 23.27 26.00 52.34 51.99 51.85 39.56
MoNE 28.16 40.24 63.12 78.00 25.21 32.40 63.33 60.65 63.54 50.52
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Avg.
0% Moonlight 58.28 82.49 80.40 92.00 67.30 45.60 81.12 65.70 71.11 71.56
25% Angular 39.76 52.69 38.90 79.00 42.57 32.20 68.50 61.01 62.04 52.96
FLAP 48.55 76.01 75.93 90.00 55.84 42.20 77.97 64.26 68.19 66.55
MC-SMoE 47.61 73.15 78.72 89.00 46.11 43.60 80.36 56.32 71.43 65.14
RS 55.80 80.64 78.69 90.00 46.73 46.40 81.01 58.84 72.30 67.82
MoNE 55.89 80.60 79.57 90.00 55.23 46.80 80.85 61.01 71.98 69.10
50% Angular 27.90 28.54 48.01 49.00 25.67 28.80 52.56 51.99 47.75 40.02
FLAP 33.87 61.36 63.30 75.00 36.80 36.00 69.37 57.04 62.12 54.98
MC-SMoE 29.52 47.94 59.54 79.00 23.94 31.40 67.30 57.04 60.46 50.68
RS 37.80 58.42 70.86 89.00 23.27 38.00 78.18 57.76 70.80 58.23
MoNE 43.09 70.03 76.12 90.00 23.57 40.80 78.78 58.84 70.17 61.27
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Avg.
0% Deepseek-V2-Lite 48.72 76.18 79.88 88.00 54.96 43.60 80.25 61.37 71.51 67.16
25% Angular 32.00 53.28 64.92 75.00 26.95 34.00 71.33 58.84 61.01 53.04
FLAP 43.69 71.46 75.26 84.00 47.28 41.40 78.18 62.82 67.72 63.53
MC-SMoE 36.69 60.77 71.31 84.00 42.22 36.60 75.57 58.48 68.67 59.37
RS 49.32 74.41 69.39 90.00 50.35 43.80 80.14 62.09 70.24 65.53
MoNE 46.67 74.62 78.47 90.00 49.05 43.00 79.76 62.09 71.43 66.12
50% Angular 24.06 32.79 40.40 61.00 23.22 26.80 56.42 57.76 49.09 41.28
FLAP 35.24 60.31 69.66 79.00 36.13 35.20 74.76 56.68 64.09 56.79
MC-SMoE 24.57 35.82 56.36 67.00 28.29 27.40 61.37 52.35 53.91 45.23
RS 36.01 57.45 57.98 89.00 24.91 40.80 78.02 54.15 62.75 55.67
MoNE 37.20 67.17 73.39 84.00 42.30 36.80 75.30 59.57 67.88 60.40
Baseline methods
We selected structured pruning methods for different structures as baseline. Notably, unless explicitly stated, we did not apply any weight update to compare the effect of pruning methods. Specifically, for general structured pruning methods, we used Angular for layer pruning [11], which evaluates the layer importance by the angular distance between the input activations for different layers, and we used FLAP for weight pruning [2], which evaluates the channel importance by the fluctuation of the input activations and compensates the performance loss with the averaged output activations. For expert pruning methods, we adopted the expert merging method in MC-SMoE [18] for one of the expert pruning baselines. Another expert pruning baseline is RS [12], which uses routing scores to evaluate the expert importance and discards less accessed ones. As there are 64 experts and 8 activated experts per layer for the three MoE models, the two expert pruning methods will not change the activated parameters. In contrast, the proposed MoNE adaptively reduces the activated parameters for tokens routed to novices.
Implementation details
We tested two pruning ratios: 25% and 50%. To demonstrate the robustness of MoNE to calibration data, we conducted experiments on two calibration data sources: Zyda2 [31] and C4 [27]. Both datasets are constructed for LLM pretraining and C4 is commonly used as the calibration dataset for model pruning [19, 9, 11, 34]. Besides, we also investigated the performance under three calibration sample sizes: 100, 500 and 1000 in Section 5.3.
To evaluate performance recovery capabilities, we conducted continued pretraining on the 25% compressed OLMoE model pruned by 100 Zyda2 samples, as only this model releases its pretraining dataset, OLMoE-mix-0924333https://huggingface.co/datasets/allenai/OLMoE-mix-0924. The sequence length was set to 4096 and the global token size per step was 4M. Each pruned model was trained with 2B tokens, i.e., 512 steps, and the peak and minimum learning rate (lr) were 5e-5 and 5e-6, respectively. We employed the cosine lr scheduler with 50 warm up steps. Other hyperparameters were the same as the original configuration for OLMoE [25]. All the experiments could run on a single H20 GPU, but we accelerated the training with 16 H20 GPUs.
Following previous researches [23, 3, 19, 34, 2], we adopted lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness [10] to measure the zero shot accuracy and average results on nine downstream tasks: Arc-c and Arc-e [6], BoolQ [5], COPA [28], MMLU [13], OBQA [24], PIQA [4], RTE [33] and Winogrande [29].
(a) 25% pruning ratio.
(b) 50% pruning ratio.
Figure 3: Average accuracy versus accuracy drop variance. MoNE advances the Pareto frontier across varying model architectures, calibration data sources and calibration sample sizes.
5.2 Effectiveness evaluation
This section validates the effectiveness of MoNE by comparing the zero shot performance with 100 calibration samples from the Zyda2 dataset. The results are presented in Table 1. This table indicates that MoNE consistently outperforms baseline methods in terms of the average accuracy on the nine tasks. In particular, compared to baseline methods, it improves average accuracy by 2.72 for the 25% pruned OLMoE and by 3.61 for the 50% pruned Deepseek-V2-Lite. Notably, it incurs accuracy drop as small as only 1.04 for the 25% pruned Deepseek-V2-Lite. Furthermore, MoNE-pruned models achieves either the best or the second best result for individual tasks only except for the 50% pruned Moonlight on the MMLU task.
An interesting observation is that all the three expert pruning methods, MC-SMoE, RS and MoNE can achieve results on par with or even better than the original models on certain tasks. The specific examples are shown in green background in Table 1. All these results indicate that there is indeed redundancy existing in the expert level for the three pretrained MoE models, and expert pruning can rule out such redundancy to achieve even better results on these tasks. Besides, compared to MC-SMoE and RS which keep the same activated parameters, MoNE introduces less activated parameters for tokens routed to novices. Nevertheless, MoNE consistently surpasses the two baseline methods, demonstrating its strong capability.
5.3 Robustness evaluation
This section evaluates the robustness of MoNE across three key dimensions: model architecture (OLMoE, Moonlight and Deepseek-V2-Lite), calibration data sources (Zyda2 and C4), and calibration sample sizes (100, 500 and 1000). For each dimension, we vary one factor while averaging results over the other two, measuring both average accuracy and the standard deviation of accuracy drop. The results are visualized in Figure 3, with detailed scores provided in Appendix A. As shown in Figure 3(a), MoNE advances the Pareto frontier across all three dimensions at the 25% pruning ratio, demonstrating superior robustness and effectiveness compared to existing structured pruning methods. At the 50% pruning ratio (Figure 3(b)), MoNE exhibits slightly higher variance under varying model architectures and calibration sample sizes. Nevertheless, it remains the most effective method, outperforming baseline methods by a significant margin of 2.85.
5.4 Ablation study
Figure 4: Ablation study on expert access frequency, output variance and novice replacement. Numbers are the difference to the proposed MoNE. The detailed result is provided in Appendix C.
This section presents the ablation study to evaluate the effects of the two redundancy metrics and the impact of novice replacement across the downstream tasks. Figure 4 displays the average accuracy drop relative to our proposed methods, with lower values indicating greater degradation. Results are averaged over the three evaluation dimensions to provide a robust assessment.
We observe that integrating the fused expert redundancy score with novice replacement yields better performance, particularly under higher pruning ratios. This indicates that our approach is especially effective in preserving model quality when pruning is more aggressive. Notably, for tasks such as BoolQ, COPA, and PIQA, our proposed method outperforms the ablation baselines by a large margin—achieving accuracy gains of up to 8.85. However, for MMLU, pruning based solely on frequency appears to offer a slight advantage, suggesting that frequently activated experts may play a more critical role in domain-specific reasoning tasks.
Table 2: Zero shot performance of the 25% pruned OLMoE after continued pretraining with 2B tokens from OLMoE-mix-0924. Best results are in bold, and the second best are underlined.
Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
OLMoE 49.23 76.89 70.09 85.00 53.54 44.40 79.76 71.84 68.90 66.63
Angular 38.82 64.69 63.52 82.00 25.42 39.80 76.50 51.62 59.04 55.71
FLAP 42.24 69.07 69.51 80.00 45.56 40.40 77.42 50.18 63.54 59.77
MC-SMoE 42.75 70.41 69.76 80.00 44.13 37.60 75.79 66.43 64.96 61.31
RS 44.97 72.94 70.73 85.00 43.28 43.00 78.67 72.20 65.98 64.09
MoNE 47.35 74.33 71.56 87.00 43.30 40.40 78.89 67.51 67.25 64.18
5.5 Accuracy recovery with continued pretraining
This section demonstrates the capability of the MoNE framework with the 25% pruned OLMoE after continued pretraining. The results are summarized in Table 2. This table shows that MoNE achieves the average accuracy closest to the original model with only 2B tokens from a pretraining dataset, demonstrating the promising capability of the MoNE computation framework. Besides, MC-SMoE and RS reclaim 8.54 and 14.02 average accuracy, indicating that expert pruning is not only effective to eliminate redundancy, but also relatively easier to recover performance with pretraining.
5.6 Inference latency and memory footprint
Figure 5: Inference latency and memory footprint over downstream tasks for the three models with different pruning ratios and methods. Latency are normalized with original models. Bars are normalized latency and lines are maximum allocated memory.
This section evaluates the inference latency and memory footprint for pruned models. For each model, we used the original HuggingFace inference implementation. Experiments were conducted on a single H20 GPU, using the first 20 samples from each downstream task with batch size 1 to eliminate padding effects. The maximum generation length was fixed to 256. Though different calibration data might affect the performance, no such variation was observed under this setup.
Figure 5 depicts the normalized latency and maximum allocated memory for models pruned by different methods and pruning ratios. Since the batch size is 1 and generation length is short, the main memory overhead is from model parameters. With the same pruning ratio, different methods achieve similar peak memory reduction. In contrast, different structured pruning methods can yield varying generation latency for MoE models. Angular prunes layers, thus achieving latency reduction linear to pruning ratio. RS and MC-SMoE show descent speedup on OLMoE model but cannot accelerate other two models. This may because the two models need to move routing scores to CPU to decide the expert execution, which downgrades the GPU utilization. Notably, MoE models pruned by FLAP cannot show any speedup. This may be attributed to the tile-quantization effect from the small weight matrices with irregular shapes within each expert [1]. MoNE shows speedup around 10% over original models, suggesting that a more efficient inference implementation is necessary to fully unlock the computational advantages of MoNE. We consider this as promising future work.
6 Conclusion
In this paper, we proposes MoNE, a novel expert pruning method that replaces redundant experts with lightweight novices to compress MoE models. MoNE evaluates the expert redundancy based on the access frequency and output variance of each expert within a transformer layer. Experts with low usage and stable output are pruned and replaced by novices, which are the unbiased estimation of their outputs to compress the model size while incurring minimal performance degradation. Extensive experiments on different MoE model architectures, calibration data sources and calibration sample sizes reveal that MoNE exceeds existing structured pruning methods by maintaining higher zero shot performance across nine downstream tasks, demonstrating the effectiveness and robustness of MoNE.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A More detailed results
Table 3: Zero shot performance with 500 calibration samples from Zyda2 dataset.
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% OLMoE 49.23 76.89 70.09 85.00 53.54 44.40 79.76 71.84 68.90 66.63
25% Angular 32.76 61.91 61.71 74.00 23.13 37.60 71.65 53.07 55.09 52.33
FLAP 37.03 63.43 64.28 81.00 41.12 38.80 72.63 54.51 63.54 57.37
MC-SMoE 33.36 54.46 70.03 81.00 37.05 33.80 68.17 64.62 65.19 56.41
RS 23.81 40.91 57.92 69.00 27.79 30.20 63.71 50.18 57.38 46.77
MoNE 41.04 65.66 70.24 87.00 41.21 40.00 76.61 64.98 66.61 61.48
50% Angular 27.22 37.50 53.91 62.00 23.96 26.60 58.27 52.35 51.85 43.74
FLAP 30.38 52.99 62.17 70.00 30.91 33.20 66.65 59.21 57.06 51.40
MC-SMoE 25.43 32.28 54.80 66.00 22.95 25.40 55.22 54.51 54.14 43.41
RS 25.34 28.24 42.48 56.00 23.07 26.40 52.34 52.35 51.78 39.78
MoNE 26.28 36.03 65.17 75.00 26.02 29.00 61.64 57.40 62.67 48.80
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Moonlight 58.28 82.49 80.40 92.00 67.30 45.60 81.12 65.70 71.11 71.56
25% Angular 39.76 52.69 38.90 79.00 42.57 32.20 68.50 61.01 62.04 52.96
FLAP 48.55 76.05 77.49 89.00 55.12 42.40 76.66 65.34 68.59 66.58
MC-SMoE 37.46 63.47 76.82 81.00 48.27 35.00 71.11 58.84 70.17 60.24
RS 55.63 79.46 78.93 91.00 46.60 45.80 80.90 59.93 72.14 67.82
MoNE 55.03 78.96 79.36 90.00 54.39 45.40 80.69 58.48 71.74 68.23
50% Angular 27.90 28.54 48.01 49.00 25.67 28.80 52.56 51.99 47.75 40.02
FLAP 34.98 61.32 65.14 73.00 37.93 35.60 69.53 56.68 62.35 55.17
MC-SMoE 22.87 29.34 58.93 72.00 23.91 26.00 55.06 54.51 52.72 43.93
RS 37.71 59.97 71.65 89.00 25.48 38.00 76.22 57.04 68.82 58.21
MoNE 38.23 64.48 75.90 87.00 23.89 39.20 77.42 58.48 70.56 59.46
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Deepseek-V2-Lite 48.72 76.18 79.88 88.00 54.96 43.60 80.25 61.37 71.51 67.16
25% Angular 32.00 53.28 64.92 75.00 26.95 34.00 71.33 58.84 61.01 53.04
FLAP 43.86 72.18 75.93 85.00 47.22 41.80 78.45 62.09 68.27 63.87
MC-SMoE 33.53 52.95 73.67 81.00 41.68 32.20 66.70 52.35 70.17 56.03
RS 48.98 73.23 71.77 89.00 52.68 44.60 79.33 61.01 70.32 65.66
MoNE 44.62 73.11 78.01 90.00 48.29 41.80 79.43 59.21 71.35 65.09
50% Angular 24.06 32.79 40.40 61.00 23.22 26.80 56.42 57.76 49.09 41.28
FLAP 33.11 60.73 67.77 78.00 31.87 36.80 72.91 55.23 63.93 55.59
MC-SMoE 25.34 32.07 47.98 59.00 26.28 25.60 56.20 54.51 53.75 42.30
RS 38.14 62.42 53.03 86.00 38.98 38.80 74.37 48.01 64.64 56.04
MoNE 36.69 66.04 73.36 86.00 41.07 35.60 75.73 58.12 69.69 60.26
Table 4: Zero shot performance with 1000 calibration samples from Zyda2 dataset.
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% OLMoE 49.23 76.89 70.09 85.00 53.54 44.40 79.76 71.84 68.90 66.63
25% Angular 32.76 61.91 61.71 74.00 23.13 37.60 71.65 53.07 55.09 52.33
FLAP 38.91 66.20 64.65 79.00 40.05 37.60 74.65 62.82 63.77 58.63
MC-SMoE 38.31 61.66 61.87 73.00 33.85 33.20 66.38 57.04 65.43 54.53
RS 26.62 43.35 59.76 70.00 27.50 30.00 65.02 49.46 56.67 47.60
MoNE 42.32 64.52 66.45 88.00 41.55 40.80 78.02 64.26 67.01 61.44
50% Angular 27.22 37.50 53.91 62.00 23.96 26.60 58.27 52.35 51.85 43.74
FLAP 30.46 54.88 62.32 70.00 29.63 32.00 67.95 57.04 57.54 51.31
MC-SMoE 25.43 35.31 55.11 67.00 22.92 25.40 54.52 51.99 51.14 43.20
RS 24.23 29.50 41.44 58.00 23.45 24.20 51.25 50.54 51.07 39.30
MoNE 26.71 37.21 65.44 75.00 26.50 31.80 63.22 55.96 63.46 49.48
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Moonlight 58.28 82.49 80.40 92.00 67.30 45.60 81.12 65.70 71.11 71.56
25% Angular 39.76 52.69 38.90 79.00 42.57 32.20 68.50 61.01 62.04 52.96
FLAP 49.49 76.52 76.91 90.00 54.47 42.00 77.26 65.70 69.06 66.82
MC-SMoE 34.81 57.74 77.22 85.00 36.13 34.60 71.49 58.12 71.35 58.50
RS 56.23 79.25 79.02 91.00 46.76 45.20 80.63 59.57 72.14 67.76
MoNE 55.55 79.29 79.66 90.00 54.79 45.60 80.69 59.21 72.14 68.55
50% Angular 27.90 28.54 48.01 49.00 25.67 28.80 52.56 51.99 47.75 40.02
FLAP 36.95 63.09 64.56 71.00 37.14 34.80 69.86 57.40 63.22 55.33
MC-SMoE 23.12 41.04 61.83 68.00 25.44 26.80 58.43 53.07 53.75 45.72
RS 39.42 64.56 70.98 90.00 24.76 40.60 77.58 58.12 68.98 59.45
MoNE 39.42 65.82 76.15 88.00 23.81 40.40 78.40 58.12 70.88 60.11
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Deepseek-V2-Lite 48.72 76.18 79.88 88.00 54.96 43.60 80.25 61.37 71.51 67.16
25% Angular 32.00 53.28 64.92 75.00 26.95 34.00 71.33 58.84 61.01 53.04
FLAP 44.28 72.39 76.12 85.00 47.16 41.20 78.29 62.82 67.40 63.85
MC-SMoE 39.33 64.31 71.53 85.00 41.84 40.00 76.12 58.48 69.53 60.68
RS 49.23 73.32 71.38 89.00 52.44 45.00 79.71 60.29 71.19 65.73
MoNE 44.71 73.99 78.35 88.00 49.19 42.20 79.27 59.93 70.88 65.17
50% Angular 24.06 32.79 40.40 61.00 23.22 26.80 56.42 57.76 49.09 41.28
FLAP 32.59 60.77 68.13 77.00 31.43 36.80 73.94 54.87 61.64 55.24
MC-SMoE 33.19 47.81 55.35 81.00 24.91 33.80 67.52 51.99 64.33 51.10
RS 37.12 61.95 54.53 88.00 39.15 38.40 75.46 51.62 62.98 56.58
MoNE 37.71 65.36 73.49 84.00 41.22 36.60 75.30 57.40 69.53 60.07
Table 5: Zero shot performance with 100 calibration samples from C4 dataset.
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% OLMoE 49.23 76.89 70.09 85.00 53.54 44.40 79.76 71.84 68.90 66.63
25% Angular 32.76 61.91 61.71 74.00 23.13 37.60 71.65 53.07 55.09 52.33
FLAP 36.01 62.67 64.83 75.00 36.31 37.00 75.73 58.84 62.75 56.57
MC-SMoE 32.76 51.05 54.71 80.00 23.05 37.80 70.89 53.43 66.61 52.26
RS 34.56 50.38 63.64 85.00 26.51 39.60 76.01 55.96 64.01 55.07
MoNE 36.69 55.18 67.03 86.00 25.53 40.20 77.80 56.68 67.64 56.97
50% Angular 27.22 37.50 53.91 62.00 23.96 26.60 58.27 52.35 51.85 43.74
FLAP 26.62 50.51 62.14 62.00 26.32 28.80 69.86 52.71 57.06 48.45
MC-SMoE 25.68 32.24 57.65 60.00 22.83 27.40 58.60 52.35 51.38 43.12
RS 23.12 31.99 45.02 59.00 23.80 24.80 53.59 51.62 49.57 40.28
MoNE 26.37 39.65 62.29 75.00 22.93 32.00 66.10 52.71 61.01 48.67
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Moonlight 58.28 82.49 80.40 92.00 67.30 45.60 81.12 65.70 71.11 71.56
25% Angular 39.76 52.69 38.90 79.00 42.57 32.20 68.50 61.01 62.04 52.96
FLAP 46.76 75.04 75.90 85.00 52.98 41.00 78.07 65.70 67.40 65.32
MC-SMoE 48.38 74.20 78.69 92.00 50.41 44.20 81.18 54.87 71.19 66.12
RS 55.97 79.76 78.93 91.00 52.74 46.20 81.39 59.21 72.38 68.62
MoNE 54.27 79.25 78.07 90.00 53.78 45.60 81.34 57.76 72.14 68.02
50% Angular 27.90 28.54 48.01 49.00 25.67 28.80 52.56 51.99 47.75 40.02
FLAP 31.14 54.63 62.69 73.00 32.26 30.40 70.40 58.12 60.93 52.62
MC-SMoE 29.10 52.23 57.22 85.00 22.92 36.40 71.71 52.71 63.54 52.31
RS 37.80 58.42 70.86 89.00 23.27 38.00 78.18 57.76 70.80 58.23
MoNE 33.87 57.07 72.75 90.00 22.97 38.80 78.62 61.01 70.17 58.36
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Deepseek-V2-Lite 48.72 76.18 79.88 88.00 54.96 43.60 80.25 61.37 71.51 67.16
25% Angular 32.00 53.28 64.92 75.00 26.95 34.00 71.33 58.84 61.01 53.04
FLAP 39.68 66.58 76.45 84.00 36.14 39.80 78.67 57.04 67.48 60.65
MC-SMoE 36.52 59.30 59.76 81.00 36.80 37.20 75.30 54.87 69.22 56.66
RS 49.32 74.41 69.39 90.00 50.35 43.80 80.14 62.09 70.24 65.53
MoNE 47.44 74.03 77.71 90.00 49.10 42.60 80.25 63.18 70.24 66.06
50% Angular 24.06 32.79 40.40 61.00 23.22 26.80 56.42 57.76 49.09 41.28
FLAP 29.95 52.36 68.72 75.00 23.45 34.20 75.46 52.71 61.09 52.55
MC-SMoE 28.75 36.57 59.45 82.00 23.65 30.60 64.96 51.62 59.75 48.60
RS 36.01 57.45 57.98 89.00 24.91 40.80 78.02 54.15 62.75 55.67
MoNE 35.24 55.47 68.20 87.00 22.94 35.40 76.50 53.43 68.35 55.84
Table 6: Zero shot performance with 500 calibration samples from C4 dataset.
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% OLMoE 49.23 76.89 70.09 85.00 53.54 44.40 79.76 71.84 68.90 66.63
25% Angular 32.76 61.91 61.71 74.00 23.13 37.60 71.65 53.07 55.09 52.33
FLAP 36.26 63.80 63.49 72.00 35.09 36.00 75.35 53.07 61.96 55.23
MC-SMoE 26.71 48.61 65.90 69.00 35.14 31.00 61.92 54.87 61.09 50.47
RS 34.56 49.41 63.85 86.00 27.23 39.80 75.57 56.32 64.40 55.24
MoNE 37.88 55.89 65.23 86.00 25.03 41.20 77.42 57.40 67.80 57.10
50% Angular 27.22 37.50 53.91 62.00 23.96 26.60 58.27 52.35 51.85 43.74
FLAP 26.71 50.42 62.20 64.00 27.52 30.20 70.29 52.35 55.09 48.75
MC-SMoE 23.98 32.07 61.77 63.00 23.01 25.40 54.79 57.40 52.49 43.77
RS 22.95 31.78 49.11 59.00 23.75 23.00 53.59 53.43 50.04 40.74
MoNE 29.01 42.51 62.29 79.00 22.98 32.00 70.84 52.71 60.06 50.16
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Moonlight 58.28 82.49 80.40 92.00 67.30 45.60 81.12 65.70 71.11 71.56
25% Angular 39.76 52.69 38.90 79.00 42.57 32.20 68.50 61.01 62.04 52.96
FLAP 45.22 73.40 72.29 84.00 53.81 40.40 78.13 64.98 67.32 64.40
MC-SMoE 48.38 74.16 78.56 89.00 48.78 43.00 81.12 58.48 71.03 65.84
RS 55.72 79.80 79.20 91.00 52.93 44.60 81.23 59.57 71.74 68.42
MoNE 54.52 78.87 78.56 90.00 53.80 45.20 81.07 57.04 71.74 67.87
50% Angular 27.90 28.54 48.01 49.00 25.67 28.80 52.56 51.99 47.75 40.02
FLAP 31.83 54.50 63.18 74.00 29.18 31.40 71.16 55.96 61.88 52.57
MC-SMoE 33.45 54.59 63.82 79.00 23.95 33.20 72.31 52.71 61.09 52.68
RS 39.59 58.67 70.49 86.00 23.36 38.60 78.02 54.15 70.24 57.68
MoNE 33.28 56.40 71.96 91.00 22.95 37.60 78.73 58.12 70.64 57.85
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Deepseek-V2-Lite 48.72 76.18 79.88 88.00 54.96 43.60 80.25 61.37 71.51 67.16
25% Angular 32.00 53.28 64.92 75.00 26.95 34.00 71.33 58.84 61.01 53.04
FLAP 40.53 64.81 77.43 84.00 37.70 39.60 79.43 57.04 67.32 60.87
MC-SMoE 39.33 61.99 60.95 84.00 41.37 38.60 77.75 58.48 62.04 58.28
RS 49.66 74.49 65.63 91.00 50.04 43.60 79.92 61.01 69.77 65.01
MoNE 46.59 73.19 77.37 89.00 48.80 43.40 79.87 62.09 70.80 65.68
50% Angular 24.06 32.79 40.40 61.00 23.22 26.80 56.42 57.76 49.09 41.28
FLAP 30.97 52.27 68.47 77.00 23.16 36.40 75.08 52.71 62.75 53.20
MC-SMoE 23.55 35.73 57.43 71.00 23.86 32.20 61.10 51.62 51.30 45.31
RS 36.95 58.71 51.16 87.00 24.23 40.20 77.64 54.15 61.72 54.64
MoNE 34.39 55.72 67.68 84.00 22.90 35.60 76.28 54.51 67.72 55.42
Table 7: Zero shot performance with 1000 calibration samples from C4 dataset.
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% OLMoE 49.23 76.89 70.09 85.00 53.54 44.40 79.76 71.84 68.90 66.63
25% Angular 32.76 61.91 61.71 74.00 23.13 37.60 71.65 53.07 55.09 52.33
FLAP 36.35 61.24 64.50 72.00 37.08 36.00 75.41 57.40 62.59 55.84
MC-SMoE 40.10 68.22 61.65 70.00 36.91 39.60 71.65 59.57 57.93 56.18
RS 34.39 49.41 64.89 85.00 26.64 40.60 75.95 57.04 63.46 55.26
MoNE 38.74 57.45 65.57 86.00 24.59 41.80 77.42 56.68 68.43 57.41
50% Angular 27.22 37.50 53.91 62.00 23.96 26.60 58.27 52.35 51.85 43.74
FLAP 26.79 49.92 62.17 63.00 27.34 31.60 70.24 53.43 56.35 48.98
MC-SMoE 26.96 41.58 58.56 56.00 23.00 28.60 60.23 52.35 50.83 44.23
RS 23.81 32.28 46.09 58.00 23.76 24.00 53.59 51.99 48.86 40.26
MoNE 29.78 42.80 62.32 80.00 22.99 33.60 71.49 52.71 60.06 50.64
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Moonlight 58.28 82.49 80.40 92.00 67.30 45.60 81.12 65.70 71.11 71.56
25% Angular 39.76 52.69 38.90 79.00 42.57 32.20 68.50 61.01 62.04 52.96
FLAP 45.22 74.20 72.78 86.00 53.40 41.20 78.35 65.34 68.90 65.04
MC-SMoE 51.62 76.52 77.86 92.00 43.08 43.60 80.25 55.60 71.11 65.74
RS 55.97 79.59 78.75 90.00 53.01 45.20 81.34 59.57 71.90 68.37
MoNE 53.75 79.38 78.53 90.00 53.75 45.80 81.39 57.76 71.59 67.99
50% Angular 27.90 28.54 48.01 49.00 25.67 28.80 52.56 51.99 47.75 40.02
FLAP 31.31 55.26 63.73 72.00 32.98 33.00 71.16 56.32 62.75 53.17
MC-SMoE 28.75 49.20 58.90 75.00 23.72 33.40 66.32 53.79 59.75 49.87
RS 38.74 58.88 69.85 88.00 23.32 38.00 77.80 61.01 70.17 58.42
MoNE 34.81 57.58 71.96 92.00 22.95 38.60 78.18 57.04 70.72 58.20
Pruning ratio Model/Method Arc-c Arc-e BoolQ COPA MMLU OBQA PIQA RTE Winogrande Average
0% Deepseek-V2-Lite 48.72 76.18 79.88 88.00 54.96 43.60 80.25 61.37 71.51 67.16
25% Angular 32.00 53.28 64.92 75.00 26.95 34.00 71.33 58.84 61.01 53.04
FLAP 39.33 66.50 76.33 84.00 38.14 40.60 78.84 59.93 67.88 61.28
MC-SMoE 37.29 58.42 65.72 81.00 37.91 39.00 76.12 55.60 67.80 57.65
RS 49.23 74.33 68.44 88.00 51.59 44.40 80.25 62.45 69.69 65.38
MoNE 45.99 72.98 77.43 89.00 48.90 42.60 79.98 61.01 71.59 65.50
50% Angular 24.06 32.79 40.40 61.00 23.22 26.80 56.42 57.76 49.09 41.28
FLAP 31.66 53.41 69.36 75.00 23.04 35.40 75.41 52.71 62.51 53.17
MC-SMoE 28.16 36.49 56.94 81.00 25.76 31.80 64.69 52.71 58.09 48.40
RS 39.08 61.32 52.54 83.00 24.58 39.40 77.86 54.51 63.46 55.08
MoNE 34.04 55.51 68.26 85.00 22.92 36.40 76.44 54.15 66.61 55.48
This section presents the experiment results on Zyda2 dataset with 500 and 1000 samples in Table 3 and Table 4. Table 5, Table 6 and Table 7 presents the experiment results on C4 dataset with 100, 500 and 1000 samples. The observations are the similar to those in Section 5.2 and 5.3.
Appendix B Redundant expert visualization
Figure 6: Layer-wise normalized expert access frequency and output variance of OLMoE for Arc-C & Arc-E, MMLU and Winogrande.
Figure 7: Layer-wise normalized expert access frequency and output variance of OLMoE for OBQA, PIQA and RTE.
Figure 8: Layer-wise normalized expert access frequency and output variance of OLMoE for BoolQ and COPA.
Figure 9: Layer-wise normalized expert access frequency and output variance of Moonlight for Arc-C & Arc-E, MMLU and Winogrande.
Figure 10: Layer-wise normalized expert access frequency and output variance of Moonlight for OBQA, PIQA and RTE.
Figure 11: Layer-wise normalized expert access frequency and output variance of Moonlight for BoolQ and COPA.
Figure 12: Layer-wise normalized expert access frequency and output variance of Deepseek-V2-Lite for OBQA, PIQA and RTE.
Figure 13: Layer-wise normalized expert access frequency and output variance of Deepseek-V2-Lite for BoolQ and COPA.
This section complements the visualization of redundant experts for different models across the downstream tasks. The results are depicted in Figure 6 - Figure 13. As mentioned in Figure 1, for each figure, expert in blue circles has both high frequency and variance. Expert in red circles only has high variance. Expert in green circles only has high frequency. For each model across the nine downstream tasks, we can always identify the same important experts, validating the effectiveness of the redundancy metric, i.e., the expert access frequency and output variance.
Appendix C Comprehensive ablation study results
Figure 14: Ablation study on expert access frequency, output variance and novice replacement with detailed results. Numbers are the difference to the proposed MoNE.
Figure 14 reports the detailed ablation study on the impacts of the three factors: expert access frequency, output variance and novice replacement. The results in this figure validates that the three factors play an important role in maintaining the effectiveness of the pruned models on different model architectures, calibration data sources and calibration sample sizes. Fusing the three factors ensures the robustness of the proposed MoNE.
Appendix D Limitations
The proposed MoNE is an expert pruning method for the structured pruning of MoE models and the experiments demonstrate excellent performance. However, there are a few limitations that are not covered in this paper. First, we do not conduct experiments on larger MoE models. The largest MoE models employed in this paper are at the 16B scale. Larger models may require more redundancy features other than the two metrics employed in this paper. However, those large models also need more computation resources for experiments. Second, we do not explore the combination of different structured pruning methods. While this paper focuses on expert pruning which is unique to MoE models and demonstrates the advantages of expert pruning over other structured pruning methods, combining different pruning methods can potentially expand the search space, leaving finer granularity to retain more model capabilities. We leave the two avenues as promising directions for future work.